# A Cross-Cultural Analysis of Social Norms in Bollywood and Hollywood Movies

### v2

@@ A.5 LIWC Correlation Results
Tables A6, A7, A8 and A9 contain the significantly correlated ($p<0.05$) LIWC categories, the most frequent five words for each category, pearson r and 95% confidence interval.
LIWC Categories Top-5 words correlation 95% CI
EMO_NEG (bad, mad, scared, worry, fear) 0.330 [0.315, 0.344]
TONE_NEG (lost, kill, wrong, bad, hit) 0.249 [0.233, 0.265]
POWER (sir, respect, own, kill, poor) 0.198 [0.182, 0.214]
EMOTION (love, good, bad, happy, crazy) 0.168 [0.152, 0.185]
YOU (you, your, you’re, yourself, you’ve) 0.161 [0.145, 0.178]
SOCREFS (you, your, he, her, him) 0.148 [0.132, 0.165]
FEELING (feel, touch, feeling, felt, hard) 0.133 [0.117, 0.150]
DRIVES (we, our, us, sir, married) 0.115 [0.098, 0.132]
SOCIAL (you, your, he, her, him) 0.110 [0.094, 0.127]
MORAL (wrong, innocent, duty, decent, excuse) 0.102 [0.085, 0.118]
AFFECT (love, good, keep, respect, well) 0.078 [0.062, 0.095]
NEGATE (not, don’t, no, aren’t, won’t) 0.065 [0.049, 0.082]
FEMALE (her, she, girl, she’s, mom) 0.060 [0.043, 0.077]
PPRON (you, i, me, your, my) 0.058 [0.041, 0.075]
FAMILY (son, married, uncle, dad, mom) 0.055 [0.039, 0.072]
PREP (to, of, in, for, on) 0.053 [0.037, 0.070]
SEXUAL (chaste, lust, sex, sexy, pimp) 0.051 [0.034, 0.067]
PRONOUN (you, i, me, your, my) 0.040 [0.023, 0.057]
AUXVERB (is, are, have, be, don’t) 0.029 [0.013, 0.046]
CONFLICT (kill, killed, accusing, killing, cruel) 0.024 [0.007, 0.041]
SWEAR (hell, bloody, idiot, damn, ass) 0.024 [0.007, 0.041]
EMO_ANGER (mad, angry, hate, cruel, argue) 0.022 [0.005, 0.039]
FOCUSPRESENT (is, are, don’t, i’m, aren’t) 0.021 [0.005, 0.038]
@@
LIWC Categories Top-5 words correlation 95% CI
DRIVES (we, our, us, sir, work) 0.122 [0.104,0.140]
DET (the, a, my, your, that) 0.101 [0.082,0.119]
ACHIEVE (work, better, win, best, try) 0.098 [0.080,0.116]
POWER (sir, own, respect, kill, power) 0.091 [0.072,0.109]
SOCREFS (you, your, he, we, our) 0.085 [0.067,0.103]
PREP (of, to, in, for, with) 0.084 [0.066,0.103]
MORAL (wrong, duty, brave, arrogant, useless) 0.075 [0.057,0.093]
CONJ (and, but, so, if, as) 0.075 [0.056,0.093]
REWARD (win, won, glory, success, successful) 0.071 [0.052,0.089]
TONE _POS (love, good, thank, well, great) 0.064 [0.046,0.082]
POLITIC (nation, army, sultan, president, dynasty) 0.064 [0.045,0.082]
WE (we, our, us, we’ll, let’s) 0.060 [0.042,0.079]
SOCIAL (you, your, he, we, our) 0.060 [0.041,0.078]
FAMILY (son, papa, married, dad, uncle) 0.059 [0.041,0.078]
AFFILIATION (we, our, us, dear, we’ll) 0.059 [0.041,0.077]
FEELING (feel, feeling, hard, felt, sense) 0.055 [0.036,0.073]
ETHNICITY (indian, indians, british, hindi, caste) 0.054 [0.036,0.072]
MALE (he, his, him, son, sir) 0.054 [0.036,0.072]
CULTURE (indian, nation, army, car, indians) 0.044 [0.026,0.062]
AFFECT (love, good, thank, well, great) 0.040 [0.021,0.058]
ARTICLE (the, a, an, tha) 0.039 [0.021,0.058]
PPRON (you, i, my, your, me) 0.036 [0.017,0.054]
PROSOCIAL (thank, please, sorry, respect, gift) 0.032 [0.014,0.051]
FUNCTION (you, the, i, of, to) 0.032 [0.013,0.050]
YOU (you, your, you’re, you’ve, you’ll) 0.030 [0.011,0.048]
THEY (they, their, them, they’re, they’ll) 0.029 [0.011,0.048]
CERTITUDE (really, real, surely, proved, actually) 0.021 [0.002,0.039]
EMO _POS (love, good, happy, happiness, smile) 0.020 [0.001,0.038]
@@
LIWC Categories Top-5 words correlation 95% CI
EMO_NEG (sick, pain, fear, bad, afraid) 0.425 [0.403, 0.446]
TONE_NEG (lost, wrong, sick, pain, poor) 0.355 [0.331, 0.377]
EMOTION (good, love, sick, pain, bad) 0.290 [0.266, 0.314]
POWER (own, sir, poor, killed, war) 0.263 [0.238, 0.287]
AFFECT (well, good, love, help, damn) 0.168 [0.142, 0.193]
DRIVES (we, us, our, work, we’re) 0.152 [0.127, 0.178]
FUNCTION (you, i, the, to, of) 0.131 [0.105, 0.157]
PPRON (you, i, me, i’m, my) 0.111 [0.085, 0.137]
MORAL (wrong, excuse, decent, honest, duty) 0.110 [0.084, 0.136]
I (i, me, i’m, my, i’ll) 0.102 [0.076, 0.128]
EMO_SAD (crying, cry, sob, lonely, sad) 0.096 [0.070, 0.122]
NEGATE (no, not, don’t, nothing, never) 0.090 [0.063, 0.116]
PREP (to, of, in, for, on) 0.089 [0.062, 0.115]
PRONOUN (you, i, that, it, me) 0.088 [0.062, 0.114]
YOU (you, your, you’re, yourself, you’ve) 0.080 [0.054, 0.106]
AUXVERB (be, i’m, is, was, have) 0.077 [0.051, 0.103]
FOCUSPAST (was, did, were, been, didn’t) 0.069 [0.043, 0.096]
SOCIAL (you, your, we, he, you’re) 0.060 [0.033, 0.086]
CONJ (and, so, but, if, when) 0.056 [0.030, 0.082]
LINGUISTIC (you, i, the, to, of) 0.056 [0.030, 0.082]
ALLNONE (no, all, nothing, never, yes) 0.055 [0.029, 0.081]
FAMILY (son, dad, baby, mom, mama) 0.054 [0.028, 0.081]
SOCREFS (you, your, we, he, you’re) 0.050 [0.024, 0.077]
EMO_ANX (fear, afraid, worry, terrified, scared) 0.042 [0.015, 0.068]
ILLNESS (sick, pain, pains, flu, sickly) 0.038 [0.012, 0.064]
FEELING (feel, felt, pain, feeling, hard) 0.035 [0.008, 0.061]
EMO_ANGER (hate, hated, mad, angry, hates) 0.033 [0.007, 0.059]
DIFFER (not, but, if, didn’t, or) 0.033 [0.007, 0.059]
DISCREP (should, can, would, can’t, want) 0.032 [0.006, 0.059]
COGNITION (no, not, all, know, but) 0.032 [0.005, 0.058]
@@
LIWC Categories Words correlation 95% CI
PPRON (you, i, i’m, me, my) 0.135 [0.113,0.158]
SOCREFS (you, your, we, he, you’re) 0.131 [0.108,0.154]
FAMILY (son, dad, baby, mom, mama) 0.121 [0.099,0.144]
CONJ (and, so, but, as, if) 0.114 [0.091,0.137]
SOCIAL (you, your, we, he, you’re) 0.102 [0.079,0.125]
FUNCTION (you, i, the, of, to) 0.100 [0.077,0.123]
I (i, i’m, me, my, i’ll) 0.095 [0.072,0.118]
YOU (you, your, you’re, you’ve, yourself) 0.091 [0.068,0.114]
MALE (he, his, him, man, son) 0.084 [0.061,0.107]
DRIVES (we, our, us, we’re, dad) 0.076 [0.053,0.099]
TONE _POS (good, well, thank, great, love) 0.072 [0.049,0.095]
PRONOUN (you, i, i’m, that, it) 0.072 [0.048,0.095]
AUXVERB (i’m, be, is, was, have) 0.071 [0.048,0.094]
EMO _POS (good, love, happy, hope, wonderful) 0.070 [0.047,0.093]
PREP (of, to, in, for, on) 0.070 [0.047,0.093]
AFFILIATION (we, our, us, we’re, dad) 0.054 [0.031,0.078]
EMOTION (good, love, happy, hope, bad) 0.052 [0.029,0.075]
ETHNICITY (american, irish, chinese, german, christian) 0.052 [0.029,0.075]
REWARD (win, won, winner, successful, earned) 0.048 [0.025,0.071]
ACHIEVE (work, better, best, trying, try) 0.045 [0.022,0.068]
POWER (sir, own, war, strong, mighty) 0.044 [0.021,0.067]
AFFECT (good, well, thank, great, love) 0.041 [0.018,0.064]
MORAL (wrong, excuse, hero, brave, dignity) 0.040 [0.017,0.064]
FEMALE (her, she, she’s, girl, ladies) 0.034 [0.011,0.057]
CULTURE (american, car, president, nation, mayor) 0.033 [0.009,0.056]
FOCUSPAST (was, did, been, were, had) 0.032 [0.009,0.055]
SHEHE (he, his, him, her, she) 0.031 [0.008,0.054]
WORK (work, job, school, deal, company) 0.028 [0.005,0.051]
@@
A.6 Clustering Results
Tables A12 and A13 contain the manually annotated Cluster Theme, the total number of samples in each cluster, and Bollywood vs Hollywood distribution. The distance was set to $5$ and the duplicates were removed. The theme and top three examples demonstrating its meaning are provided in Tables A10 and A11.
@@ A.6 Clustering Results
Table A12: Distribution of reasons (shame) across manually labeled clusters. A total of twenty-six clusters were generated with distance=5. Duplicates were removed for clustering. One cluster had generic reasons ( indicating lack of shame) which was removed for further analysis. Finally, 25 clusters were considered.
Table A13: Distribution of reasons (pride) across manually labeled clusters. A total of fifteen clusters were generated with distance=5. Duplicates were removed for clustering. One cluster had generic statements (without explicit reason) and was removed for further analysis. ’Achievement’ and ’Winning’ were merged to form ’Achievement’ and, ’Bravery’ and ’Nation’ were merged to form ’Bravery & Nation’. Finally, 12 clusters were considered for analysis.
(a) Bollywood
(b) Hollywood
Figure 5: Gender-wise differences in themes associated with normative expectations from shame-related discourse in Bollywood and Hollywood. $\Delta(Female,Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq. 1 but for dimension gender). Promiscuity is most associated with the female gender in both industries.
(a) Bollywood
(b) Hollywood
Figure 6: Gender-wise differences in themes associated with normative expectations from pride-related discourse in Bollywood and Hollywood. $\Delta(Female,Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq. 1 but for dimension gender). Promiscuity is most associated with the female gender in both industries.

### v3

@@ A.6 LIWC Correlation Results
Tables A7, A8, A9 and A10 contain the significantly correlated ($p<0.05$) LIWC categories, the most frequent five words for each category, pearson r and 95% confidence interval.
LIWC Categories Top-5 words r p-value 95% CI
Negative emotions (EMO_NEG) (bad, mad, scared, worry, fear) 0.330 0.000 [0.315, 0.344]
Negative tone (TONE_NEG) (lost, kill, wrong, bad, hit) 0.249 0.000 [0.233, 0.265]
POWER (sir, respect, own, kill, poor) 0.198 0.000 [0.182, 0.214]
EMOTION (love, good, bad, happy, crazy) 0.168 0.000 [0.152, 0.185]
YOU (you, your, you’re, yourself, you’ve) 0.161 0.000 [0.145, 0.178]
Social references (SOCREFS) (you, your, he, her, him) 0.148 0.000 [0.132, 0.165]
FEELING (feel, touch, feeling, felt, hard) 0.133 0.000 [0.117, 0.150]
DRIVES (we, our, us, sir, married) 0.115 0.000 [0.098, 0.132]
SOCIAL (you, your, he, her, him) 0.110 0.000 [0.094, 0.127]
MORAL (wrong, innocent, duty, decent, excuse) 0.102 0.000 [0.085, 0.118]
AFFECT (love, good, keep, respect, well) 0.078 0.000 [0.062, 0.095]
NEGATE (not, don’t, no, aren’t, won’t) 0.065 0.000 [0.049, 0.082]
FEMALE (her, she, girl, she’s, mom) 0.060 0.000 [0.043, 0.077]
Personal Pronouns (PPRON) (you, i, me, your, my) 0.058 0.000 [0.041, 0.075]
FAMILY (son, married, uncle, dad, mom) 0.055 0.000 [0.039, 0.072]
Preposition (PREP) (to, of, in, for, on) 0.053 0.000 [0.037, 0.070]
SEXUAL (chaste, lust, sex, sexy, pimp) 0.051 0.000 [0.034, 0.067]
PRONOUN (you, i, me, your, my) 0.040 0.000 [0.023, 0.057]
AUXVERB (is, are, have, be, don’t) 0.029 0.001 [0.013, 0.046]
CONFLICT (kill, killed, accusing, killing, cruel) 0.024 0.007 [0.007, 0.041]
SWEAR (hell, bloody, idiot, damn, ass) 0.024 0.007 [0.007, 0.041]
Anger (EMO_ANGER) (mad, angry, hate, cruel, argue) 0.022 0.015 [0.005, 0.039]
FOCUSPRESENT (is, are, don’t, i’m, aren’t) 0.021 0.018 [0.005, 0.038]
@@
LIWC Categories Top-5 words r p-value 95% CI
DRIVES (we, our, us, sir, work) 0.122 0.000 [0.104,0.140]
Determiner (DET) (the, a, my, your, that) 0.101 0.000 [0.082,0.119]
ACHIEVE (work, better, win, best, try) 0.098 0.000 [0.080,0.116]
POWER (sir, own, respect, kill, power) 0.091 0.000 [0.072,0.109]
Social References (SOCREFS) (you, your, he, we, our) 0.085 0.000 [0.067,0.103]
Preposition (PREP) (of, to, in, for, with) 0.084 0.000 [0.066,0.103]
MORAL (wrong, duty, brave, arrogant, useless) 0.075 0.000 [0.057,0.093]
Conjunction (CONJ) (and, but, so, if, as) 0.075 0.000 [0.056,0.093]
REWARD (win, won, glory, success, successful) 0.071 0.000 [0.052,0.089]
Positive tone (TONE_POS) (love, good, thank, well, great) 0.064 0.000 [0.046,0.082]
POLITIC (nation, army, sultan, president, dynasty) 0.064 0.000 [0.045,0.082]
WE (we, our, us, we’ll, let’s) 0.060 0.000 [0.042,0.079]
SOCIAL (you, your, he, we, our) 0.060 0.000 [0.041,0.078]
FAMILY (son, papa, married, dad, uncle) 0.059 0.000 [0.041,0.078]
AFFILIATION (we, our, us, dear, we’ll) 0.059 0.000 [0.041,0.077]
FEELING (feel, feeling, hard, felt, sense) 0.055 0.000 [0.036,0.073]
ETHNICITY (indian, indians, british, hindi, caste) 0.054 0.000 [0.036,0.072]
MALE (he, his, him, son, sir) 0.054 0.000 [0.036,0.072]
CULTURE (indian, nation, army, car, indians) 0.044 0.000 [0.026,0.062]
AFFECT (love, good, thank, well, great) 0.040 0.000 [0.021,0.058]
ARTICLE (the, a, an, that) 0.039 0.000 [0.021,0.058]
Personal Pronouns (PPRON) (you, i, my, your, me) 0.036 0.000 [0.017,0.054]
PROSOCIAL (thank, please, sorry, respect, gift) 0.032 0.001 [0.014,0.051]
FUNCTION (you, the, i, of, to) 0.032 0.001 [0.013,0.050]
YOU (you, your, you’re, you’ve, you’ll) 0.030 0.003 [0.011,0.048]
THEY (they, their, them, they’re, they’ll) 0.029 0.003 [0.011,0.048]
CERTITUDE (really, real, surely, proved, actually) 0.021 0.041 [0.002,0.039]
Positive Emotion (EMO_POS) (love, good, happy, happiness, smile) 0.020 0.049 [0.001,0.038]
@@
LIWC Categories Top-5 words r p 95% CI
Negative emotions (EMO_NEG) (sick, pain, fear, bad, afraid) 0.425 0.000 [0.403, 0.446]
Negative tone (TONE_NEG) (lost, wrong, sick, pain, poor) 0.355 0.000 [0.331, 0.377]
EMOTION (good, love, sick, pain, bad) 0.290 0.000 [0.266, 0.314]
POWER (own, sir, poor, killed, war) 0.263 0.000 [0.238, 0.287]
AFFECT (well, good, love, help, damn) 0.168 0.000 [0.142, 0.193]
DRIVES (we, us, our, work, we’re) 0.152 0.000 [0.127, 0.178]
FUNCTION (you, i, the, to, of) 0.131 0.000 [0.105, 0.157]
Personal Pronouns (PPRON) (you, i, me, i’m, my) 0.111 0.000 [0.085, 0.137]
MORAL (wrong, excuse, decent, honest, duty) 0.110 0.000 [0.084, 0.136]
I (i, me, i’m, my, i’ll) 0.102 0.000 [0.076, 0.128]
Sadness (EMO_SAD) (crying, cry, sob, lonely, sad) 0.096 0.000 [0.070, 0.122]
NEGATE (no, not, don’t, nothing, never) 0.090 0.000 [0.063, 0.116]
Preposition (PREP) (to, of, in, for, on) 0.089 0.000 [0.062, 0.115]
PRONOUN (you, i, that, it, me) 0.088 0.000 [0.062, 0.114]
YOU (you, your, you’re, yourself, you’ve) 0.080 0.000 [0.054, 0.106]
Auxiliary verb (AUXVERB) (be, i’m, is, was, have) 0.077 0.000 [0.051, 0.103]
FOCUSPAST (was, did, were, been, didn’t) 0.069 0.000 [0.043, 0.096]
SOCIAL (you, your, we, he, you’re) 0.060 0.000 [0.033, 0.086]
Conjunction (CONJ) (and, so, but, if, when) 0.056 0.000 [0.030, 0.082]
LINGUISTIC (you, i, the, to, of) 0.056 0.000 [0.030, 0.082]
ALLNONE (no, all, nothing, never, yes) 0.055 0.000 [0.029, 0.081]
FAMILY (son, dad, baby, mom, mama) 0.054 0.000 [0.028, 0.081]
Social references (SOCREFS) (you, your, we, he, you’re) 0.050 0.000 [0.024, 0.077]
Anxiety (EMO_ANX) (fear, afraid, worry, terrified, scared) 0.042 0.004 [0.015, 0.068]
ILLNESS (sick, pain, pains, flu, sickly) 0.038 0.010 [0.012, 0.064]
FEELING (feel, felt, pain, feeling, hard) 0.035 0.020 [0.008, 0.061]
Anger (EMO_ANGER) (hate, hated, mad, angry, hates) 0.033 0.026 [0.007, 0.059]
DIFFER (not, but, if, didn’t, or) 0.033 0.027 [0.007, 0.059]
Discrepancy (DISCREP) (should, can, would, can’t, want) 0.032 0.029 [0.006, 0.059]
COGNITION (no, not, all, know, but) 0.032 0.033 [0.005, 0.058]
@@
LIWC Categories Top-5 words r p-value 95% CI
Personal Pronouns (PPRON) (you, i, i’m, me, my) 0.135 0.000 [0.113,0.158]
Social References (SOCREFS) (you, your, we, he, you’re) 0.131 0.000 [0.108,0.154]
FAMILY (son, dad, baby, mom, mama) 0.121 0.000 [0.099,0.144]
Conjunction (CONJ) (and, so, but, as, if) 0.114 0.000 [0.091,0.137]
SOCIAL (you, your, we, he, you’re) 0.102 0.000 [0.079,0.125]
FUNCTION (you, i, the, of, to) 0.100 0.000 [0.077,0.123]
I (i, i’m, me, my, i’ll) 0.095 0.000 [0.072,0.118]
YOU (you, your, you’re, you’ve, yourself) 0.091 0.000 [0.068,0.114]
MALE (he, his, him, man, son) 0.084 0.000 [0.061,0.107]
DRIVES (we, our, us, we’re, dad) 0.076 0.000 [0.053,0.099]
Positive tone (TONE_POS) (good, well, thank, great, love) 0.072 0.000 [0.049,0.095]
PRONOUN (you, i, i’m, that, it) 0.072 0.000 [0.048,0.095]
Auxiliary Verb (AUXVERB) (i’m, be, is, was, have) 0.071 0.000 [0.048,0.094]
Positive Emotion (EMO_POS) (good, love, happy, hope, wonderful) 0.070 0.000 [0.047,0.093]
Preposition (PREP) (of, to, in, for, on) 0.070 0.000 [0.047,0.093]
AFFILIATION (we, our, us, we’re, dad) 0.054 0.000 [0.031,0.078]
EMOTION (good, love, happy, hope, bad) 0.052 0.000 [0.029,0.075]
ETHNICITY (american, irish, chinese, german, christian) 0.052 0.000 [0.029,0.075]
REWARD (win, won, winner, successful, earned) 0.048 0.000 [0.025,0.071]
ACHIEVE (work, better, best, trying, try) 0.045 0.000 [0.022,0.068]
POWER (sir, own, war, strong, mighty) 0.044 0.001 [0.021,0.067]
AFFECT (good, well, thank, great, love) 0.041 0.001 [0.018,0.064]
MORAL (wrong, excuse, hero, brave, dignity) 0.040 0.002 [0.017,0.064]
FEMALE (her, she, she’s, girl, ladies) 0.034 0.009 [0.011,0.057]
CULTURE (american, car, president, nation, mayor) 0.033 0.012 [0.009,0.056]
FOCUSPAST (was, did, been, were, had) 0.032 0.013 [0.009,0.055]
SHEHE (he, his, him, her, she) 0.031 0.016 [0.008,0.054]
WORK (work, job, school, deal, company) 0.028 0.033 [0.005,0.051]
@@
A.7 Clustering Results
Tables A13 and A14 contain the manually annotated Cluster Theme, the total number of samples in each cluster, and Bollywood vs Hollywood distribution. The distance was set to $5$ and the duplicates were removed. The theme and top three examples demonstrating its meaning are provided in Tables A11 and A12.
Theme Examples from cluster
Stealing shamelessness and taking money, lying and misusing money, Stealing and bribery
Poverty not having money, being poor, living in poor conditions
Incompetence not fulfilling responsibilities, not meeting expectations or making a mistake, Failure or perceived incompetence
Promiscuity being promiscuous, Having a love affair, shameless behavior and expressing love
Inappropriate behavior not living up to someone’s expectations, being disloyal, Being rude and ungrateful
Disobedience not upholding cultural values, disobedience or lack of respect, being disloyal
Parent-related not standing up for his mother, not taking care of his father, not taking responsibility for his son
Marriage-related refusing to marry, being forced into marriage, not accepting the proposed marriage
Immodesty Wearing inappropriate clothes, inappropriate behavior nudity in public, behaving indecently in public
Gender Roles not behaving as per her husband’s expectations, behaving inappropriately in front of her daughter, going against her husband
Sexual Harrassment assaulting a girl, physical assault on women, sexual assault
Cowardice lack of courage, cowardice and inability to stand up for oneself, lack of pride and integrity
Alcholism excessive drinking, inappropriate behavior due to alcohol, drinking and irresponsible behavior
Non-conformity not conforming to gender norms, breaking gender norms, behaving inappropriately according to societal norms
Illegal activities engaging in criminal activities, Being a criminal, committing illegal acts
Betrayal Betrayal or dishonesty, betrayal and infidelity, Deception/Betrayal
Lack of accountability not taking responsibility and blaming others, wrongdoing without remorse, not acknowledging wrongdoing
Disrespect disrespecting others, disrespecting an authority figure, Disrespectful and inappropriate behavior
Harm causing harm to others, committing harmful deeds, causing trouble and endangering others
Lying/Deception lying and hiding information, Lying or deceit, Deception/Not being truthful
Inappropriate sexual behavior inappropriate advances and comments, inappropriate behavior towards a young girl, inappropriate language and behavior
Social Etiquette Being humiliated in public, being disrespected and belittled, being mocked and treated shamefully
Privacy-related invading personal space, invading someone’s privacy, intrusion of personal space, trespassing/invading personal space
Accusation accused of wrongdoing, being accused of something, being accused of infidelity
Family norms Disrespecting family, causing harm and shame to family, causing difficulty and shame to family
@@ A.7 Clustering Results
Theme Examples from cluster
Achievement being a brilliant student, working hard and achieving a spectacular result, professional achievement
Family Roles raising their child, having an unselfish mother, having a simple and good daughter
Self-identity being a man, being a woman, being gay
Duty fulfilling duty, Accomplishing something great, saving the world
Bravery & Nation bravery and sacrifice, bravery and service to nation, bravery and selflessness
Doing the "right" thing for doing the right thing, for being heroes, for having courage
Son’s success son’s success and progress, son’s hard work and achievement, son’s determination
Family Honor maintaining dignity and pride, taking care of family pride, saving family’s pride
Justice standing up for justice, fighting for justice, bringing justice
Marriage-related marrying with pride and respect, sister’s marriage, daughter’s marriage into a reputed family
Appearance & Assets his manhood, his wealth and beautiful wife, his knowledge and power
Resilience resilience and determination, overcoming struggles, enduring hardships without complaint
Ethnolinguistic Identity being Indian, defending the pride of Rajputs, being a Maharashtrian
@@
Theme (Shame-related norms) Total Samples Hollywood Bollywood
Accountability 802 185 617
Accusation 70 0 70
Alcoholism 98 18 80
Betrayal 262 26 236
Cowardice 209 51 158
Disobedience 216 49 167
Disrespect 246 26 220
Family norms 203 24 179
Gender roles 221 17 204
Harm 225 64 161
Illegal activities 253 27 226
Immodesty 311 58 253
Incompetence 342 83 259
Lying/Deception 222 53 169
Marriage-related 179 25 154
Non-conformity 208 28 180
Parent-related 263 34 229
Poverty 414 108 306
Privacy-related 102 14 88
Promiscuity 595 89 506
Sexual behavior 407 33 374
Sexual Harassment 232 24 208
Social Etiquette 417 97 320
Stealing 351 49 302
Total 6848 1182 5666
Table A13: Distribution of reasons (shame) across manually labeled clusters. A total of twenty-six clusters were generated with distance=5. Duplicates were removed for clustering. One cluster with generic reasons ( such as phrases "lack of shame") was removed and two clusters with similar reasons (related to accountability) were merged. Finally, 24 clusters were considered. The total is slightly more than numbers in Table A3 since a norm could be mapped to multiple clusters.
Theme (Pride-related norms) Total Samples Hollywood Bollywood
Achievement 332 174 158
Bravery 248 69 179
Doing the "right" thing 82 82 0
Duty 792 449 343
Ethnolinguistic Identity 148 7 141
Family Honor 263 77 186
Family Roles 356 156 200
Justice 255 109 146
Nation 216 57 159
Physical Appearance 106 49 57
Resilience 180 68 112
Self-identity 273 157 116
Son’s Achievements 315 87 228
Wedding 120 23 97
Winning 411 205 206
Total 4097 1769 2328
Table A14: Distribution of reasons (pride) across manually labeled clusters. A total of fifteen clusters were generated with distance=5. Duplicates were removed for clustering. ’Achievement’ and ’Winning’ were merged to form ’Achievement’ and, ’Bravery’ and ’Nation’ were merged to form ’Bravery & Nation’. Finally, 13 clusters were considered for analysis.
(a) Shame-related norms in Bollywood
(b) Shame-related norms in Hollywood
Figure A1: Gender-wise differences in themes associated with normative expectations from shame-related discourse in Bollywood and Hollywood. $\Delta(Female,Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq. 1 but for dimension gender). * indicates insignificant difference i.e., $p>0.05$
(a) Shame-related norms in Bollywood
(b) Shame-related norms in Hollywood
Figure A2: Gender-wise differences in themes associated with normative expectations from pride-related discourse in Bollywood and Hollywood. $\Delta(Female,Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq. 1 but for dimension gender). * indicates insignificant difference i.e., $p>0.05$